OpenStudy (anonymous):

Find the area of the region that lies inside both circles r=sin(theta) and r=cos(theta

OpenStudy (anonymous):

is this in polar coordinates?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

\[A=2\int\limits_{0}^{\pi/4}1/2\sin^2\theta d \theta=\int\limits_{0}^{\pi/4}1/2(1-\cos2\theta)d \theta\] \[=1/2[\theta-1/2\sin2\theta]_{0}^{\pi/4}={1 \over 8} \pi - {1 \over 4}\]

OpenStudy (anonymous):

@AnwarA: How'd you get that answer, if I may ask?

OpenStudy (anonymous):

From what I've gathered, we can see that the point of intersection is at (theta)=π/4. Then we can get one integral of 1/2 * sin^2(theta) from 0 to π/4, added to an integral of 1/2 * cos^2(theta) from π/4 to π/2. Wouldn't that be enough to account for both halves of the petal-like enclosed shape?

OpenStudy (anonymous):

Alright, so we've gotten exactly the same answers...xD But can you show me how they're the same? I haven't done polar curves yet and I'm curious.

OpenStudy (anonymous):

well you're integrating twice over the same region sinx from 0 to pi/4

OpenStudy (anonymous):

the same area*

OpenStudy (anonymous):

that's why I multiplied the integral by 2.. make sense?

OpenStudy (anonymous):

Yep, I see now. Thanks. :)

OpenStudy (anonymous):

no problem :)

OpenStudy (anonymous):

you haven't done polar curves? you still in high school?

OpenStudy (anonymous):

or maybe a freshman?

OpenStudy (anonymous):

I'm a junior in AP Calc, polar curves is what we're doing immediately after we finish sequences & series. I skipped pre-calc, so now I'm pretty much going through the rest of the syllabus on my own to cover anything I missed.

OpenStudy (anonymous):

I see. wish you the best!

OpenStudy (anonymous):

how did you find the point of intersection?

OpenStudy (anonymous):

Plot the two polar curves (or set the two equations equal to each other) and based on simple trigonometry, the angle at which they both equal each other is π/4. I think you'll see exactly what I mean when you graph it though.

OpenStudy (anonymous):

alright thank you, it makes sense. btw im review sequences and series right now for my calculus exam can you reccemend a video or website which covers the section well? thank you

OpenStudy (anonymous):

Here you have links of tons of different types of convergence tests, basic concepts; basically anything you could want to see about these. :) No problem, http://tutorial.math.lamar.edu/Classes/CalcII/SeriesIntro.aspx